OpenStudy (anonymous):

how do you Rationalize the denominator of 4 over quantity of 6 minus square root of 7

OpenStudy (goformit100):

@uri Help

OpenStudy (psymon):

Well, you need to multiply both the top and bottom by the conjugate of the denominator. The conjugate is just taking the denominator and changing the sign of it. So this is what you would be solving for: \[\frac{ 4 }{ (6-\sqrt{7}) }*\frac{ (6+\sqrt{7}) }{{ (6+\sqrt{7}) } }\]

OpenStudy (anonymous):

so then id do something like foil?

OpenStudy (psymon):

Right.

OpenStudy (anonymous):

okay thankyou :)

OpenStudy (psymon):

Mhm. See what you get and post it if you wanna check your answer :3

OpenStudy (anonymous):

alright :)

OpenStudy (phi):

notice the bottom is in the form (a-b)(a+b) = a^2 + b^2 (it is called the "conjugate" ). the bottom turns into just a nice number

OpenStudy (psymon):

a^2 - b^2

OpenStudy (anonymous):

I got a answer but its nothing like what my teacher showed us before

OpenStudy (psymon):

Whatd ya come up with?

OpenStudy (anonymous):

4 squareroot 7 +53

OpenStudy (psymon):

Yeah, somehow that didnt turn out. I'll work it out.

OpenStudy (psymon):

\[\frac{ 4 }{ 6-\sqrt{7} }*\frac{ 6+\sqrt{7} }{ 6+\sqrt{7} }\] Starting with the numerator \[4(6+\sqrt{7}) = 24 + 4\sqrt{7}\] Now the denominator I'll work out each step at a time for. \[6*6 = 36\] \[6*\sqrt{7} = 6\sqrt{7}\] \[-\sqrt{7}*6 = -6\sqrt{7}\] \[-\sqrt{7}*\sqrt{7} = -7\] From here you think you could fit the rest of the pieces together? :P

OpenStudy (anonymous):

let me see what I come up with

OpenStudy (anonymous):

okay I got it :) thankyou so much!

OpenStudy (psymon):

Yep, np ^_^

OpenStudy (psymon):

What was mentioned before is a quick way to try and multiply conjugates, though. If you have (a+b)(a-b), you can just find a^2 - b^2 and that would be the answer. So for your problem, I could have just done 6^2 minus sqrt(7)^2 and got the same thing

OpenStudy (phi):

fix: notice the bottom is in the form (a-b)(a+b) = a^2 - b^2 (it is called the "conjugate" ). the bottom turns into just a nice number